OpenStudy (anonymous):

If two automobiles have the same velocity do they have the same acceleration?

OpenStudy (raffle_snaffle):

yes, they would have the same velocity...

OpenStudy (raffle_snaffle):

Do you know why though?

OpenStudy (anonymous):

No

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\left[ \frac{ Dv }{ Dt }\right] _{t = t1} = \left[ \frac{ Dv }{ Dt }\right] _{t = t2} aunque V(t1) \neq V(t2)\]

OpenStudy (shane_b):

Just because they have the same velocity does NOT mean they have the same acceleration.

OpenStudy (raffle_snaffle):

Think about it this way. Take two objects with different masses, drop them off a tall structure. The acc. will be the same, but the velocity is going to be different.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Acceleration is the derivative of velocity.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (shane_b):

@raffle_snaffle said: "Take two objects with different masses, drop them off a tall structure. The acc. will be the same, but the velocity is going to be different." That's not correct :) They will have different KE and PE values due to their differences in mass but their velocities will be the same throughout the fall. No where in the equations for constant acceleration will you find mass.